OpenStudy (shamil98):

An old man had a bag of rice on the first day, each day afterwards he would get triple of what he got before, but every 5 days he would receive no bags of rice. How many bags of rice would the old man have after 45 days?

OpenStudy (abb0t):

89?

OpenStudy (shamil98):

no -_-

OpenStudy (shamil98):

S = (first term)(1-r^n)/(1-r) is your formula, but you must remember 9 of those days he did not receive any rice.

OpenStudy (anonymous):

brb as I go look at wios post

OpenStudy (shamil98):

Lol batman xD

OpenStudy (anonymous):

:P

OpenStudy (anonymous):

more rice than we can fit in the USA

OpenStudy (shamil98):

LOL you shouldve seen wio's question lmao

OpenStudy (shamil98):

like 18 quintillion pounds of rice or something

OpenStudy (anonymous):

so put * 4/5 in there somewhere

OpenStudy (anonymous):

you ever verify the 6*12 piece of paper origami crease project?

OpenStudy (shamil98):

I haven't.

OpenStudy (anonymous):



OpenStudy (anonymous):

3^(666)+1=the answer, i couldn't calculate it cuz it shows overflow for the calculator

OpenStudy (shamil98):

@wio

OpenStudy (anonymous):

The accumulative amount, or just the amount received on day 45?

OpenStudy (anonymous):

We have \[ a_n = a_1r^{n-1}=1\times 3^{n-1}=3^{n-1} \]

OpenStudy (anonymous):

Where \(d\) is the number of days, we can calculate \(n\) to be: \[ n=d-\left\lfloor \frac d5 \right\rfloor \]

OpenStudy (anonymous):

so if \(d=45\) then \[ n=45-\left\lfloor \frac {45}5 \right\rfloor=45-\left\lfloor 9 \right\rfloor=36 \]

OpenStudy (shamil98):

The sum,

OpenStudy (shamil98):

My bad I should've put that in the question

OpenStudy (anonymous):

So the number of bags received on that day is: \[ a_{36}=3^{36-1}=50031545098999707\approx 5.00\times 10^{16} \]The accumulated bags would be: \[ S_n = a\frac{1-r^n}{1-r}=(1)\frac{1-(3)^{36}}{1-(3)}=75047317648499560\approx 7.50\times 10^{16} \]

OpenStudy (shamil98):

WIO YOU HAVE 99 SS CONGRATS ! :O

OpenStudy (anonymous):

part B of the question. If each bag of rice was 20" long by 12" wide with a volume of 7 gallons, how high could you stack the bags end to end from california to new york if the distance was 3250 miles?

OpenStudy (shamil98):

LOL

OpenStudy (anonymous):

231cubic inches per gallon

OpenStudy (shamil98):

63360 inches in a mile..

OpenStudy (shamil98):

if you stacked the bags it would go beyond the earth and into space

OpenStudy (shamil98):

space begins at 62 miles of altitude o.o

OpenStudy (anonymous):

bag height = 6.7375" thick, (on an engineers drafting board, we wont count sag)

OpenStudy (anonymous):

consider: we dont know how much rice is in each of those 5.00×10^16 bags, but the number of bags is = to about 1/2 the # of atoms in the universe